son 16 years younger than father.In 3 years sons age will 1\3 of his father.Find present age of sons
Found 2 solutions by alanc, MathTherapy:
Answer by alanc(27)

son 16 years younger than father.In 3 years sons age will 1\3 of his father.Find present age of sons
-------------------------------------
We see there are 2 unknown values here.
Let s = the age of the son.
Let f = the age of the father.
The first sentence translates to : s = f - 16
The second sentence translates to (s + 3) = (1/3) * (f + 3)
We have 2 equations in 2 unknowns. We should be able to solve for s and f.
s = f - 16
(s + 3) = (1/3) * (f + 3)
s + 16 = f
3 ( s + 3) = f + 3
substitute for f.
3 (s + 3) = (s + 16) + 3
3s + 9 = s + 19
3s - s + 9 = 19
2s + 9 = 19
2s + 9 - 9 = 19 - 9
2s = 10
s = 10/2
s = 5
Therefore f = s + 16 = 5 + 16 = 21
The father is 21 years old and the son is 5 years old.

Answer by MathTherapy(10552)


son 16 years younger than father.In 3 years sons age will 1\3 of his father.Find present age of sons
Let son's age be S
Then father's age is: S + 16
We then get the following equation: S + 16 + 3 = 3(S + 3)
Solve this for S, the son's current age.
That's all. No COMPLEX and/or CONFUSING calculations.